OpenStudy (anonymous):

the bus headed north at 40 mph at 10 a.m. at noon the orange blossom special went south from the same station at 70 mph. what time was it when the train and bus were 960 miles apart.

OpenStudy (anonymous):

combine velocities = 110 mph 110 = 960 /t t = 960/110 = 8.727 hours = 8 hours 44 miniues (to nearest miniute) time was 18.44 or 6.44 pm

OpenStudy (anonymous):

ok with that?

OpenStudy (anonymous):

yes. however ive never heard of combining the velocities. but i guess it works

OpenStudy (anonymous):

yes it works when velocities are in the same direction - either going apart or coming together

OpenStudy (anonymous):

oh sorry - i made one mistake - i didnt notice the 'noon' - i thought thay both left at same time so my answer is wrong

OpenStudy (anonymous):

for the first two hours the 40mph bus travels 80 miles so the distance travelled when both of them were travelling was 960-80 = 880 so its 110 - 880 / t t = 8 hours this ia 8 hours after noon so required time is 10.00 + 2 + 8 = 20.00 hrs or 8 pm

OpenStudy (anonymous):

does the t stand for time?

OpenStudy (anonymous):

yes